# AdaSVD: Adaptive Singular Value Decomposition for Large Language Models

### v1

@@ 4.2 Main Results
We evaluate the overall performance of AdaSVD from two aspects: (1) performance under different compression ratios (40%, 50%, 60%, 70%, and 80%), (2) performance on different LLMs. Some generated contents by the compressed LLMs are provided in the supplementary file to provide a more straightforward comparison.
Performance under Different Compression Ratios. First, we evaluate the performance of LLaMA2-7B compressed by AdaSVD, vanilla SVD and the SOTA method SVD-LLM (Wang et al., 2024b) under compression ratios ranging from 40% to 80% on all datasets, as shown in Table 1. On the three language modeling datasets, AdaSVD consistently outperforms vanilla SVD, and SVD-LLM across all the compression ratios. More importantly, AdaSVD exhibits significant advantages over the baselines under higher compression ratios. These results indicate that AdaSVD is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices, which often have limited memory and processing capabilities. On the seven common sense reasoning datasets, AdaSVD also maintains its edge and performs better than the best-performing baseline on most of the datasets and consistently achieves higher average accuracy across all the compression ratios.
Performance on Different LLMs. To demonstrate the generability of AdaSVD across different LLMs, we compare AdaSVD and the baselines on four different models OPT-6.7B, LLaMA 2-7B, Vicuna-7B, and Mistral-7B – under 60% compression ratio on WikiText-2. As shown in Table 2, AdaSVD consistently outperforms vanilla SVD, FWSVD, ASVD and SVD-LLM on all LLMs, and exhibits more stable performance across different LLMs, especially compared to vanilla SVD and FWSVD. We reproduce FWSVD, ASVD, and SVD-LLM using their official GitHub repositories. FWSVD and ASVD may fail on certain LLMs with compression ratios under 60%, whereas SVD-LLM and AdaSVD maintain reasonable perplexity in such cases.
| Method | OPT-6.7B | LLaMA 2-7B | Mistral-7B | Vicuna-7B |
| SVD | 118,600 | 65,186 | 30,378 | 78,705 |
| FWSVD | nan | 62,502 | 5,481 | nan |
| ASVD | 10,326 | nan | 22,706 | nan |
| SVD-LLM | 92.10 | 89.90 | 72.17 | 64.06 |
| AdaSVD | 86.64 | 60.08 | 67.22 | 64.06 |
@@ 4.3 Ablation Study
We provide extensive ablation study results in Table 3d to show the effect of some key components in our work.
@@
| AdaSVD | 60% | 0.45 | 60.08 | 2,137.28 | 294.26 |
| AdaSVD | 60% | 0.50 | 78.82 | 6,929.39 | 339.31 |
Effectiveness of Adaptive Compensation. To validate the effectiveness of the proposed adaComp, we compare the PPL results of Llama2-7B with and without adaComp on Wikitest-2, PTB, and C4 datasets in Table 3a. Results of 70% and 80% compression ratios can be found in the supplementary file. It can be observed that AdaSVD consistently outperforms SVD-LLM after applying adaComp, and the performance gap is more significant under high compression ratios (i.e., 60%, 70%, and 80%).
Iteration Number. To investigate the impact of the number of adaComp iterations under different compression ratios, we perform an ablation study with 1, 3, and 15 iterations, as shown in Table 3c. Results for 70% and 80% compression ratios are provided in the supplementary file. At lower compression ratios (e.g., 40%, 50%, and 60%), it is observed that just 1 iteration of adaComp already outperforms the state-of-the-art method, SVD-LLM. However, increasing the number of iterations may lead to overfitting due to the limited calibration data, resulting in a performance drop. In contrast, at higher compression ratios (e.g., 70% and 80%), additional iterations lead to performance improvements, indicating that AdaSVD is more effective in high compression ratio scenarios where previous methods still struggle. This highlights the importance of balancing the number of iterations with the available data to avoid over-optimization, especially in low compression scales.
Effectiveness of Adaptive Compression Ratio. To validate the effectiveness of our adaCR, we compared the results after removing adaCR (i.e., using constant compression ratios for all layers) from AdaSVD. As shown in Table 3b, AdaSVD already outperforms SOTA SVD-LLM without using adaCR, while integrating adaCR can further enhance the performance across all compression ratios.
Minimum Retention Ratio. The minimum retention ratio () in adaCR is also crucial, and we investigate the impact of different values in Table 3d for 40%, 50%, and 60% compression ratios. It can be observed that remains relatively robust at lower compression ratios (40% and 50%), but requires more careful consideration at higher compression ratios (60%).
@@ 4.4 Integrate with Weight Quantization
Similar to previous SVD-based compression methods (Hsu et al., 2022a; Yuan et al., 2024; Wang et al., 2024b), our AdaSVD is orthogonal to other types of compression techniques. Following Wang et al. (2024b), we integrate AdaSVD with the widely used weight quantization method GPTQ (Frantar et al., 2023). As shown in Table 4, we compare AdaSVD with SVD-LLM (Wang et al., 2024b) on the LLaMA2-7B model, using different compression ratios (40%, 50%, 60%, 70%, and 80%) across the WikiText-2, PTB, and C4 datasets. The results demonstrate that, when combined with the 4-bit weight quantization method GPTQ, AdaSVD also consistently outperforms SOTA baseline SVD-LLM across all compression ratios.
@@ 5 Conclusion
In this work, we propose AdaSVD, an adaptive SVD-based compression method for LLMs. AdaSVD first proposes adaComp, which adaptively compensates for the error caused by the truncation of singular matrices, efficiently reducing compression error without requiring additional training. Furthermore, AdaSVD proposes adaCR, which adaptively assigns compression ratios based on the importance of each layer, further enhancing performance while maintaining the same target compression rate. Both strategies effectively minimize SVD compression errors, particularly at high compression ratios. Our experiments on multiple open-source LLM families demonstrate that AdaSVD pushes the performance boundary beyond the current state-of-the-art SVD-based LLM compression methods.

### v3

@@ 4.2 Main Results
We evaluate the overall performance of AdaSVD from three aspects: (1) performance under different compression ratios (40%, 50%, 60%, 70%, and 80%), (2) performance on different LLMs. (3) performance on visual language models. Some performance evaluation results and generated contents by the compressed LLMs are included in the supplementary file to provide a more straightforward comparison.
Performance under Different Compression Ratios. First, we evaluate the performance of LLaMA2-7B compressed by AdaSVD, vanilla SVD and the SOTA method SVD-LLM [40] under compression ratios ranging from 40% to 80% on all datasets, as shown in Tab. 1. On the three language modeling datasets, AdaSVD consistently outperforms vanilla SVD, and SVD-LLM across all the compression ratios. More importantly, AdaSVD exhibits significant advantages over the baselines under higher compression ratios. These results indicate that AdaSVD is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices, which often have limited memory and processing capabilities. On the five common sense reasoning datasets, AdaSVD also maintains its edge and performs better than the best-performing baseline on most of the datasets and consistently achieves higher average accuracy across all the compression ratios. Due to page limitations, comparisons for 70% and 80% compression ratios are provided in the supplementary file.
Performance on Different LLMs. To demonstrate the generability of AdaSVD across different LLMs, we compare AdaSVD and the baselines on four different models OPT-6.7B, LLaMA2-7B, Vicuna-7B, and Mistral-7B – under 60% compression ratio on WikiText-2. As shown in Tab. 2, AdaSVD consistently outperforms vanilla SVD, FWSVD, ASVD and SVD-LLM on all LLMs, and exhibits more stable performance across different LLMs, especially compared to vanilla SVD and FWSVD. We reproduce FWSVD, ASVD, and SVD-LLM using their official GitHub repositories. FWSVD and ASVD fail on these LLMs with compression ratios under 60%, whereas SVD-LLM and AdaSVD maintain reasonable perplexity in such cases.
Performance on Visual Language Models. Note that our AdaSVD can also be applied to visual language models (VLMs) like LLaVA [23]. Following Lin et al. [22], we apply SVD compression to the language part of the VLMs since it dominates the model size. As shown in Figure 5, AdaSVD shows better image captioning results than vanilla SVD and SVD-LLM on COCO dataset [3] under 40% compression ratio. More image captioning comparisons with various compression ratios can be found in supplementary file.
@@ 4.3 Ablation Study
We provide extensive ablation study results in Tab. 3 to show the effect of some key components in our work.
@@
AdaSVD | 60% | 0.35 | 53.17 | 1,608.19 | 256.66 |
AdaSVD | 60% | 0.40 | 60.08 | 2,137.29 | 294.26 |
Effectiveness of Adaptive Compensation. To validate the effectiveness of the proposed adaComp, we compare the PPL results of Llama2-7B with and without adaComp on Wikitest-2, PTB, and C4 datasets in LABEL:tab:adacomp. Results of 70% and 80% compression ratios can be found in the supplementary file. It can be observed that AdaSVD consistently outperforms SVD-LLM after applying adaComp, and the performance gap is more significant under high compression ratios (i.e., 60%, 70%, and 80%).
Iteration Number. To investigate the impact of the number of adaComp iterations under different compression ratios, we perform an ablation study with 1, 3, and 15 iterations, as shown in LABEL:tab:num_iter. Results for 70% and 80% compression ratios are provided in the supplementary file. At lower compression ratios (e.g., 40%, 50%, and 60%), it is observed that just 1 iteration of adaComp already outperforms the state-of-the-art method, SVD-LLM. However, increasing the number of iterations may lead to overfitting due to the limited calibration data, resulting in a performance drop. In contrast, at higher compression ratios (e.g., 70% and 80%), additional iterations lead to performance improvements, indicating that AdaSVD is more effective in high compression ratio scenarios where previous methods still struggle. This highlights the importance of balancing the number of iterations with the available data to avoid over-optimization, especially in low compression scales.
@@
Effectiveness of Adaptive Compression Ratio. To validate the effectiveness of our adaCR, we compared the results after removing adaCR (i.e., using constant compression ratios for all layers) from AdaSVD. As shown in LABEL:tab:adacr, AdaSVD already outperforms SOTA SVD-LLM without using adaCR, while integrating adaCR can further enhance the performance across all compression ratios.
Minimum Retention Ratio. The minimum retention ratio () in adaCR is also crucial, and we investigate the impact of different values in LABEL:tab:mrr for 40%, 50%, and 60% compression ratios (70% and 80% in supplementary file). It can be observed that remains relatively robust at lower compression ratios (40% and 50%), while contributing more at higher compression ratios (60%).
@@ 4.4 Integrate with Weight Quantization
Similar to previous SVD-based compression methods [17, 42, 40], our AdaSVD is orthogonal to other types of compression techniques. Following Wang et al. [40], we integrate AdaSVD with the widely used weight quantization method GPTQ [11]. As shown in Tab. 4, we compare AdaSVD with SVD-LLM [40] on the LLaMA2-7B model, using different compression ratios (40%, 50%, 60%, 70%, and 80%) across the WikiText-2, PTB, and C4 datasets. The results demonstrate that, when combined with the 4-bit weight quantization method GPTQ, AdaSVD also consistently outperforms SOTA SVD-LLM across all compression ratios. Under high compression ratios (i.e., 60%, 70%, and 80%), AdaSVD + GPTQ-INT4 even surpasses SVD-LLM.
@@ 5 Conclusion
In this work, we propose AdaSVD, an adaptive SVD-based compression method for LLMs. AdaSVD first proposes adaComp, which adaptively compensates for the error caused by the truncation of singular matrices, efficiently reducing compression error without requiring additional training. Furthermore, AdaSVD proposes adaCR, which adaptively assigns compression ratios based on the importance of each layer, further enhancing performance while maintaining the same target compression rate. Both strategies effectively minimize SVD compression errors, particularly at high compression ratios. Our experiments on multiple open-source LLM and VLM families demonstrate that AdaSVD pushes the performance boundary beyond the current state-of-the-art SVD-based LLM compression methods.